niyyliyah:

There are 4 grams of fiber in 1/2 cup of oats. How many grams of fiber are in 3 1/2 cups of oats?

Alexis1415:

Are there any answer choices?

niyyliyah:

answer choices : A. 28 B. 12 C. 14 D. 7

niyyliyah:

Which expression is equivalent to 4 (3n - 5)? A:7n - 20 B:12n -5 C:12n - 20 D:7n - 9

umm:

So with what we have so far: \[4=\frac{ 1 }{ 2 }\] And we already know \[\frac{ 1 }{ 2 }+\frac{ 1 }{ 2 }=1\] Count how me halves you can make 3 into, so: \[3 \times 2= 6\] Adding in the extra half of the oats\[6+1=7\] So if 4 grams is already a half, we'll go based off of how many cups of oats we have: \[7 \times 4 =28\] So it's be 28 grams of fiber.